### v1

@@ Abstract
Keywords:
Program Synthesis Language Models Syntax-Guided Synthesis
@@ 1 Introduction
In this paper, we set out to investigate whether off-the-shelf large language models can solve formal program synthesis problems. We carefully craft a library of prompts, which enables us to solve roughly $50\%$ of the SyGuS competition benchmarks. We hypothesize that, in the cases where the LLM returns only incorrect solutions, the correct solutions are most often in the vicinity of the incorrect solutions, and that, by searching in the neighborhood of the incorrect solutions, we may be able to guide an enumerative synthesize to find a solution faster. To that end, we construct a probabilistic Context-Free Grammar (pCFG) based on the incorrect solutions proposed by the LLM, and use this to guide an enumerative synthesizer within a CounterExample Guided Inductive Synthesis (CEGIS) loop.
Our final contribution is a full integration of these two techniques in a novel CEGIS algorithm with an inline syntactic oracle, in the form of an LLM that is queried by an enumerative synthesis phase. We incorporate information obtained during the synthesis search into the queries, prompting the LLM with partially enumerated functions, incorrect solutions, and counterexamples, and requesting that it provide “helper functions”, which we use to update the pCFG guiding the enumerator.
@@ 2 Background
A context-free grammar is a 4-tuple $G=(V,\Sigma,R,S)$. $V$ is a finite set of variables also known as non-terminal symbols. $\Sigma$ with $\Sigma\cap V=\emptyset$ is called the set of terminal symbols or alphabet. $R\subseteq V\times(V\cup\Sigma)\operatorname*{\!{}^{\ast}}$ is a finite relation describing the production rules of the grammar. We define $R_{\Sigma}=R\cap V\times\Sigma\operatorname*{\!{}^{\ast}}$, i.e. the set of rules restricted to those whose right-hand side only consists of terminal symbols. Elements of $(V\cup\Sigma)\operatorname*{\!{}^{\ast}}$ are known as words in sentential form. $S\in V$ is called the start symbol of the grammar $G$.
@@
In general, program synthesis is concerned with the generation (i.e., synthesis) of a program that satisfies a certain specification. Syntax-guided synthesis (SyGuS) describes a standardized function synthesis format that precisely defines a synthesis problem within first-order theories [7]. We will use the notation $\phi[F\mapsto f]$ denotes the replacing of all occurrences of $F$ in $\phi$ with $f$ while substituting all arguments to $f$ by the arguments of $F$ in the same order.
@@
SyGuS closely follows the syntax and semantics of SMT, and hence $T$ usually refers to theories that are also common in SMT. Usually, SMT solvers are queried in the background of SyGuS solvers to verify solution candidates. This connection is made explicit in Counter-Example Guided Inductive Synthesis (CEGIS) [38]. CEGIS describes enumeration search algorithms where an enumerator enumerates solution candidates with the help of a set of counter-examples and a verifier is queried to check enumerated solution candidates and produce counter-examples in case of failure. An example SyGuS problem is shown in Example 3.
@@ 3 Overview
Figure 1: An overview of pCFG-synth. Both the verifier and the LLM have access to the specification $\phi$ (in the form of the prompt for the LLM).
@@
Figure 2: An overview of iLLM-synth. Both the verifier and the enumerator have access to the specification $\phi$.
@@ 4 Stand-alone LLM
In this section, we describe how we prompt the LLM as a stand-alone synthesizer. These prompting techniques are then also deployed by pCFG-synth. We use GPT-3.5-turbo as the LLM. We rename any functions and variables in the SyGuS benchmarks to generic names to avoid the LLM producing solutions solely based on the function names.
@@ 4.1 Prompting the LLM
Prompting is an art rather than a science, but we hypothesize that it is better to ask the LLM to give a solution in a language that is more common in the training data, and then request it to translate it into our desired SMT-LIB, and experiment with both Python and Lisp. We observe that, on a subset of 50 benchmarks, we observed that soliciting responses in Lisp resulted in a 6% enhancement in the resolution of benchmarks compared to utilizing Python (and both were better than asking for SMT-lib directly). It is also reported in the literature that the efficacy of LLMs can be augmented by including emotional stimuli within prompts [27]. Incorporating the emotional prompt leads to an 8% increase in the likelihood of generating accurate solutions compared to standard, non-emotional prompts in our study. Details of our initial prompting experiments are in Appendix 0.A.
@@
Prompt 0: Prompt for Example 3.
@@
Response 0: Response from LLM after Prompt 4.1.
@@
Prompt 0: Prompt requesting a revised solution.
@@ 4.1.1 Prompts for invariant synthesis
Invariant synthesis is a specific instance of program synthesis: given a pre-condition $pre(x)$, transition-relation $trans(x,x^{\prime})$ and post-condition $post(x)$, the synthesizer is required to provide an invariant $inv$ that satisfies the following constraint: $\forall x,x^{\prime}.pre(x)\implies inv(x)\wedge(inv(x)\wedge trans(x,x))\implies inv(x^{\prime})\wedge inv(x)\implies post(x)$. We find that LLMs appear to struggle to reason about constraints presented in the above format. Inspired by “chain-of-thought” [41] prompting, where the LLM is asked to provide a step-by-step explanation, we augment our prompting strategy for invariants by asking the LLM first to explain the constraints. After requesting this explanation, we follow the same interactive prompt strategy as before.
@@ 4.1.2 Lisp to SMT-LIB Converter
Prompt 0: Request for converting Lisp to SMT-LIB code for response 4.1.
@@
Program 0: LLM-Generated program for Example 3.
@@ 5 Synthesis with pCFG Guidance: pCFG-synth
We hypothesize that, if the LLM did not propose a correct solution, the correct solution is likely to be roughly in the same “area” as the incorrect solutions it suggested, and so our synthesis algorithm aims to prioritize this area when searching for candidate programs. For simplicity, we use a simple weighted Context-Free Grammar to represent the area of solutions proposed by the LLM. We then present methods for searching the space: the first is a probabilistic top-down search, shown in Algorithm 2; the second is based on an adaptation of the $A^{*}$ algorithm [18, 26], and we integrate both into CEGIS searches as shown in Algorithm 1.
@@ 5.1 Inferring a Weighted CFG
Given a context-free grammar $G$, and a sentence $s$, the sentence is in the language of the grammar if $S\rightarrow\operatorname*{\!{}^{\ast}}s$, where $S$ is the start symbol of the grammar. The derivation of $s$ from $S$ is a sequence of rules such that $S\xrightarrow{r_{0}}s_{1}\xrightarrow{r_{1}}\ldots s_{n}\xrightarrow{r_{n}}s$ and $r_{0}\ldots r_{n}\in R$. We denote the derivation of $s$ by the sequence of rules $r_{0},\ldots r_{n}$ as $D_{s}=\{r_{0},\ldots r_{n}\}$. The left-most derivation is a derivation such that all rules expand the left-most nonterminal symbol in the sentential form.
@@
where $|r_{i}|$ is the number of times $r_{i}$ appears in the derivation. For example, consider Response 4.1.2: the weights are calculated as $w[r_{1}]=3,w[r_{2}]=3,w[r_{3}]=3,w[r_{4}]=4,w[r_{5}]=3$. These correspond to the rules from Example 3:
@@ 5.2 Probabilistic Guided Search
Algorithm 2 pCFG-synth
1:procedure pCFG-synth($prompts,\phi,G$)
2:     $conv\leftarrow[\,\,]$
3:     $progs\leftarrow\emptyset$
4:     while $prompts\neq\emptyset$ do
5:         $response\leftarrow\textsc{LLM}(prompts.pop(),conv)$
6:         $conv.append(response)$
7:         $currentProg\leftarrow\textsc{ExtractProgram}(response)$
8:         if $\forall\vec{x}\,\phi(currentProg,\vec{x})$ then
9:              return $currentProg$
10:         else
11:              $progs\leftarrow progs\cup currentProg$
12:     $W\leftarrow\textsc{WeightCounter}(prog,G)$
13:     $W_{G}\leftarrow(G,W)$
14:     $prog\leftarrow\textsc{CEGIS}(W_{G},\phi)$
15:     return $prog$
The aim of our algorithm is thus to search the area of programs closest to those with the highest weights in the wCFG, or highest probabilities in the corresponding pCFG. We devise and implement two search methods for doing this: the first is a probabilistic top-down search. To this end, we first introduce the notion of a grammar tree.
@@
To implement our probabilistic guided search, we extend this definition to a probabilistic grammar tree. Given a pCFG, $P_{G}=(V,\Sigma,R,S,\mathbb{P})$, a probabilistic grammar tree $\mathcal{T}(P_{G})$ is a directed labeled graph as defined before, but each edge has a corresponding weight $\omega$ given by $\mathbb{P}$. We limit the edges to only those needed for the left-most derivations, and so $\mathcal{E}$ and $\omega$ are defined as follows:
@@
27:procedure CompleteProgram($prog,P_{G}$$\triangleright$ Replaces nonterminal symbols with terminal symbols
@@
30:         $(\alpha\times\beta)\sim Cat(|\pi[\alpha]|,\{\mathbb{P}_{\Sigma}[\pi[\alpha]_{1}],\mathbb{P}_{\Sigma}[\pi[\alpha]_{2}],\ldots\})$triangleright Sample from distribution
@@
We search this grammar tree using a top-down enumerative synthesizer, shown in Algorithm 3. This enumerates possible programs in the grammar in a top-down manner, expanding non-terminals by randomly sampling from the categorical distribution over the production rules. That is, the search algorithm starts by considering the node corresponding to the start symbol $S$. It then chooses the next node by sampling from a categorical distribution with event probabilities corresponding to the probabilities on the outgoing edges of the current node. The categorical distribution is a generalization of the Bernoulli distribution and describes the possible results of a random variable that can take one of $K$ possible categories, with the probability of each category separately specified. Formally, to sample a rule $\alpha\times\beta$ to apply to a nonterminal symbol $\alpha$, we sample from the distribution:
@@
We then apply the sampled rule, and repeat the process. We use $prog.\{\alpha\rightarrow\beta\}$ to indicate the result of substituting the first occurence of $\alpha$ in a partial program $prog$ with $\beta$.
@@ 5.3 Weighted $A^{*}$ Search
We implement a second variation of both pCFG-synth using the $A^{*}$ weighted search algorithm as the underlying enumerator. $A^{*}$ is a search algorithm that chooses which paths to extend based on minimizing the cost of the path so far and an estimate of the cost required to extend the path to the goal, i.e., it expands nodes that minimizes $f(x)=c(x)+g(x)$, where $c(x)$ is the cost of the path to $x$ so far and $g(x)$ is the estimated cost of reaching a goal node from $x$. This technique was first used for guiding synthesis by Lee et al. [26], and we adapted the algorithm from their work.
@@
7:         for $nt\in(prog\times nt^{\prime})$ donormal-$\triangleright$ For all applicable rules
8:              $prog\leftarrow prog.\{nt\rightarrow nt^{\prime}\}$triangleright apply rule to $prog$
9:              $Q\leftarrow Q\cup(c(prog^{\prime})+g(prog^{\prime}),prog^{\prime})$
@@
The function $g(x)$ can then be thought of as the product of the probability of each nonterminal symbol in $x$ being converted into a terminal symbol.
@@ 6.1 Integrated Prompting
We construct a prompt that asks the LLM to provide helper functions to assist a student in writing SMT-lib code. We give the LLM the constraints from the target synthesis problem and the partially complete program at the point the enumerator calls the LLM. If the LLM fails to solve the problem with this prompt, we later add the most recently failed candidate solution and the counterexample it failed on. These prompts are shorter than the prompts in those used in Section 4 and, therefore, cheaper and faster to run. An example Prompt 6.1 is as follows:
@@
Prompt 0: Integrated prompt for Example 3.
@@ 6.2 Updating the Weighted Grammar
We initialize our algorithm with a weight of $1$ for each rule in the grammar. We use the LLM-generated helper functions to augment the grammar in the following way: first, any helper functions will be added directly as new production rules to replace non-terminals of the correct type in the grammar. That is, if the LLM proposes the defined function $f$, a set of rules of the form $V_{i}\times f$ are added to the grammar, for all nonterminal symbols $V_{i}$ such that this rule results in syntactically correct expressions, i.e., $V_{i}$ must be of the same type as the co-domain of $f$. This is sufficient to guarantee syntactically correct expressions because any functions proposed by the LLM that are otherwise not well formed, e.g., they reference variables that are not defined, are discarded. Any new rules are given a weight equal to the average of all the current weights for rules relevant to that non-terminal. The response parser also updates the weights of all existing rules in the grammar, according to Equation 1, calculated from the set of helper functions the LLM proposed.

### v2

@@ 1 Introduction
In this paper, we set out to investigate whether off-the-shelf large language models can solve formal program synthesis problems. We craft a library of prompts, which enables us to solve roughly $50\%$ of the SyGuS competition benchmarks. We hypothesize that, in the cases where the LLM returns only incorrect solutions, the correct solutions are most often in the vicinity of the incorrect solutions, and that, by searching in the neighborhood of the incorrect solutions, we may be able to guide an enumerative synthesizer to find a solution faster. To that end, we construct a probabilistic Context-Free Grammar (pCFG) based on the incorrect solutions proposed by the LLM, and use this to guide an enumerative synthesizer within a CounterExample Guided Inductive Synthesis (CEGIS) loop.
Our final contribution is a full integration of these techniques in a novel CEGIS algorithm with an inline syntactic oracle, in the form of an LLM that is queried by an enumerative synthesis phase. We incorporate information obtained during the synthesis search into the queries, prompting the LLM with partially enumerated functions, incorrect solutions, and counterexamples, and requesting that it provide “helper functions”, which we use to update the pCFG guiding the enumerator.
@@ 2 Background
A context-free grammar is a 4-tuple $G=(V,\Sigma,R,S)$. $V$ is a finite set of variables also known as non-terminal symbols. $\Sigma$ with $\Sigma\cap V=\emptyset$ is called the set of terminal symbols or alphabet. $R\subseteq V\times(V\cup\Sigma)\operatorname*{\!{}^{\ast}}$ is a finite relation describing the production rules of the grammar. We define $R_{\Sigma}=R\cap V\times\Sigma\operatorname*{\!{}^{\ast}}$, i.e. the set of rules restricted to those whose right-hand side only consists of terminal symbols. Elements of $(V\cup\Sigma)\operatorname*{\!{}^{\ast}}$ are known as words in sentential form. $S\in V$ is the start symbol of the grammar $G$.
@@
In general, program synthesis is concerned with the generation (i.e., synthesis) of a program that satisfies a certain specification. Syntax-guided synthesis (SyGuS) describes a standardized function synthesis format that precisely defines a synthesis problem within first-order theories [8]. We will use the notation $\phi[F\mapsto f]$ to denote the replacing of all occurrences of $F$ in $\phi$ with $f$ while substituting all arguments to $f$ by the arguments of $F$ in the same order.
@@
SyGuS closely follows the syntax and semantics of SMT, and hence $T$ usually refers to theories that are also common in SMT. Usually, SMT solvers are queried in the background of SyGuS solvers to verify solution candidates. This connection is made explicit in Counter-Example Guided Inductive Synthesis (CEGIS) [39]. CEGIS is a family of algorithms that alternate between a synthesis phase, which searches for a candidate solution that works for a subset of inputs, and a verification phase, where the candidate is checked against all possible inputs. If the verification fails, a counterexample is passed back to the synthesis phase and appended to the subset of inputs used to guide the search. The synthesis phase is often implemented as an enumerative search. An example SyGuS problem is shown in Example 1.
@@ 3 Overview
Figure 1: An overview of pCFG-synth. Both the verifier and the LLM have access to the specification $\phi$ (which is used to generate the prompt for the LLM, as well as to check whether candidate programs are correct).
@@
Figure 2: An overview of iLLM-synth. Both the verifier and the enumerator have access to the specification $\phi$ (which is used to generate the prompt for the LLM, as well as to check whether candidate programs are correct)
@@ 4 Stand-alone LLM
In this section, we describe how we prompt the LLM as a stand-alone synthesizer. These prompting techniques are then also deployed by pCFG-synth. We use GPT-3.5-turbo as the LLM. Note that the model is not fine-tuned to this problem setting. Furthermore, we rename any functions and variables in the SyGuS benchmarks to generic names to avoid the LLM producing solutions solely based on the function names.
@@ 4.1 Prompting the LLM
Prompting is an art rather than a science, but we hypothesize that it is better to ask the LLM to give a solution in a language that is more common in the training data, and then request it to translate it into our desired SMT-LIB, and experiment with both Python and Lisp. On a subset of 50 benchmarks, we observed that soliciting responses in Lisp resulted in a 6% enhancement in the resolution of benchmarks compared to utilizing Python (and both were better than asking for SMT-lib directly). It is also reported in the literature that the efficacy of LLMs can be augmented by including emotional stimuli within prompts [28]. Incorporating the emotional prompt leads to an 8% increase in the likelihood of generating accurate solutions compared to standard, non-emotional prompts in our study. Details of our initial prompting experiments are in the extended version of this paper  [^1] https://arxiv.org/html/2403.03997.
@@
[…]
Prompt 1: Prompt for Example 1.
@@
[…]
Response 1: Response from LLM after Prompt 1.
@@
[…]
Prompt 2: Prompt requesting a revised solution.
@@ 4.1.1 Prompts for invariant synthesis
Invariant synthesis is a specific instance of program synthesis: given a pre-condition $pre(x)$, transition-relation $trans(x,x^{\prime})$ and post-condition $post(x)$, the synthesizer is required to provide an invariant $inv$ that satisfies the following constraint: $\forall x,x^{\prime}.pre(x)\implies inv(x)\,\,\wedge(inv(x)\wedge trans(x,x))\implies inv(x^{\prime})\,\,\wedge inv(x)\implies post(x).$ We find that LLMs struggle to reason about constraints presented in the above format. Inspired by “chain-of-thought” [42] prompting, where the LLM is asked to provide a step-by-step explanation, we augment our prompting strategy for invariants by asking the LLM first to explain the constraints. After requesting this explanation, we follow the same interactive prompt strategy as before.
@@ 4.1.2 Lisp to SMT-LIB Converter
[…]
Prompt 3: Request for converting Lisp to SMT-LIB code for response 1.
@@
[…]
Program 1: LLM-Generated program for Example 1.
@@ 5 Synthesis with pCFG Guidance: pCFG-synth
We hypothesize that, if the LLM did not propose a correct solution, the correct solution is likely to be roughly in the same “area” as the incorrect solutions it suggested, and so our synthesis algorithm aims to prioritize this area when searching for candidate programs. For simplicity, we use a simple weighted Context-Free Grammar to represent the area of solutions proposed by the LLM. We then present methods for searching the space: the first is a probabilistic top-down search, shown in Algorithm 3; the second is based on an adaptation of the $A^{*}$ algorithm [19, 27], and we integrate both into CEGIS searches as shown in Algorithm 1. The verification phase in Algorithm 1 is implemented via a call to an SMT solver, which checks, for a candidate solution $f$, whether there exists an input such that the specification is violated, i.e., $\exists x.\neg\phi[F\mapsto f]$.
@@ 5.1 Inferring a Weighted CFG
Given a context-free grammar $G$, and a sentence $s$, the sentence is in the language of the grammar if $S\rightarrow\operatorname*{\!{}^{\ast}}s$, where $S$ is the start symbol of the grammar. The derivation of $s$ from $S$ is a sequence of rules such that $S\xrightarrow{r_{0}}s_{1}\xrightarrow{r_{1}}\ldots s_{n}\xrightarrow{r_{n}}s$ and $r_{0}\ldots r_{n}\in R$. We denote the derivation of $s$ by the sequence of rules $r_{0},\ldots r_{n}$ as $D_{s}=\{r_{0},\ldots r_{n}\}$. The left-most derivation is a derivation such that all rules expand the left-most non-terminal symbol in the sentential form.
@@
where $|r_{i}|$ is the number of times $r_{i}$ appears in the derivation. For example, consider Response 1: the weights are calculated as $w[r_{1}]=3,w[r_{2}]=3,w[r_{3}]=3,w[r_{4}]=4,w[r_{5}]=3$. These correspond to the rules from Example 1:
@@ 5.2 Probabilistic Guided Search
The aim of our algorithm is thus to search the area of programs closest to those with the highest weights in the wCFG, or highest probabilities in the corresponding pCFG. We adapt and implement two search methods for doing this: the first is a probabilistic top-down search. To this end, we first introduce the notion of a grammar tree.
@@
To implement our probabilistic guided search, we extend this definition to a probabilistic grammar tree. Given a pCFG, $P_{G}=(V,\Sigma,R,S,\mathbb{P})$, a probabilistic grammar tree $\mathcal{T}(P_{G})$ is a directed labelled graph as defined before, but each edge has a corresponding weight $\omega$ given by $\mathbb{P}$. We limit the edges to only those needed for the left-most derivations, and so $\mathcal{E}$ and $\omega$ are defined as follows:
@@
27:procedure CompleteProgram($prog,P_{G}$$\triangleright$ Replaces non-terminal symbols with terminal symbols
@@
30:         $(\alpha\times\beta)\sim Cat(|\pi[\alpha]|,\{\mathbb{P}_{\Sigma}[\pi[\alpha]_{1}],\mathbb{P}_{\Sigma}[\pi[\alpha]_{2}],\ldots\})$triangleright Sample
@@
We search this grammar tree using a top-down enumerative synthesizer, shown in Algorithm 2. This enumerates possible programs in the grammar in a top-down manner, expanding non-terminals by randomly sampling from the categorical distribution over the production rules. That is, the search algorithm starts by considering the node corresponding to the start symbol $S$. It then chooses the next node by sampling from a categorical distribution with event probabilities corresponding to the probabilities on the outgoing edges of the current node. The categorical distribution is a generalization of the Bernoulli distribution and describes the possible results of a random variable that can take one of $K$ possible categories, with the probability of each category separately specified. Formally, to sample a rule $\alpha\times\beta$ to apply to a non-terminal symbol $\alpha$, we sample from the distribution:
@@
We then apply the sampled rule, and repeat the process. We use $prog.\{\alpha\rightarrow\beta\}$ to indicate the result of substituting the first occurrence of $\alpha$ in a partial program $prog$ with $\beta$.
@@
Algorithm 3 pCFG-synth
1:procedure pCFG-synth($prompts,\phi,G$)
2:     $conv\leftarrow[\,\,]$
3:     $progs\leftarrow\emptyset$
4:     while $prompts\neq\emptyset$ do
5:         $response\leftarrow\textsc{LLM}(prompts.pop(),conv)$
6:         $conv.append(response)$
7:         $currentProg\leftarrow\textsc{ExtractProgram}(response)$
8:         if $\forall\vec{x}\,\phi(currentProg,\vec{x})$ then
9:              return $currentProg$
10:         else
11:              $progs\leftarrow progs\cup currentProg$
12:     $W\leftarrow\textsc{WeightCounter}(prog,G)$
13:     $W_{G}\leftarrow(G,W)$
14:     $prog\leftarrow\textsc{CEGIS}(W_{G},\phi)$
15:     return $prog$
@@ 5.3 Weighted $A^{*}$ Search
We implement a second variation of pCFG-synth using the $A^{*}$ weighted search algorithm as the underlying enumerator. $A^{*}$ is a search algorithm that chooses which paths to extend based on minimizing the cost of the path so far and an estimate of the cost required to extend the path to the goal, i.e., it expands nodes that minimizes $f(x)=c(x)+g(x)$, where $c(x)$ is the cost of the path to $x$ so far and $g(x)$ is the estimated cost of reaching a goal node from $x$. This technique was first used for guiding synthesis by Lee et al. [27], and we adapted the algorithm from their work.
@@
7:         for $(nt\in prog)\times nt^{\prime}$ do
8:              if $(nt\times nt^{\prime})\in P_{G}.R$ then $\triangleright$ For all applicable rules
9:                  $prog\leftarrow prog.\{nt\rightarrow nt^{\prime}\}$triangleright apply rule to $prog$
10:                  $Q\leftarrow Q\cup(c(prog)+g(prog),prog)$
@@
The function $g(x)$ can then be thought of as the product of the probability of each non-terminal symbol in $x$ being converted into a terminal symbol.
@@ 6.1 Integrated Prompting
We construct a prompt that asks the LLM to provide helper functions to assist a student in writing SMT-lib code. We give the LLM the constraints from the target synthesis problem and the partially complete program at the point the enumerator calls the LLM. If the LLM fails to solve the problem with this prompt, we later add the most recently failed candidate solution and the counterexample it failed on. These prompts are shorter than the prompts in those used in Section 4 and, therefore, cheaper and faster to run. An example Prompt 4 is as follows:
@@
[…]
Prompt 4: Integrated prompt for Example 1.
@@ 6.2 Updating the Weighted Grammar
We initialize our algorithm with a weight of $1$ for each rule in the grammar. We use the LLM-generated helper functions to augment the grammar in the following way: first, any helper functions will be added directly as new production rules to replace non-terminals of the correct type in the grammar. That is, if the LLM proposes the defined function $f$, a set of rules of the form $V_{i}\times f$ are added to the grammar, for all non-terminal symbols $V_{i}$ such that this rule results in syntactically correct expressions, i.e., $V_{i}$ must be of the same type as the co-domain of $f$. This is sufficient to guarantee syntactically correct expressions because any functions proposed by the LLM that are otherwise not well-formed, e.g., they reference variables that are not defined, are discarded. Any new rules are given a weight equal to the average of all the current weights for rules relevant to that non-terminal. The response parser also updates the weights of all existing rules in the grammar, according to Equation 1, calculated from the set of helper functions the LLM proposed.